3. (60 points) In the commodity market Qd- 200-2P, Qs-2P-50. Government impose a subsidy S 10 per unit of goods. Determine: I. Find the elasticity demand and supply in the point of equilibrium. 2.How to changes consumer and produce surpluses. 3. Pw 40, what happens with consumer and producer surpluses, if the country decide to import goods 4. Government wants to restrict import and impose a tariff 10. How to change consumer and producer surpluses, government revenue, DWL
1
Expert's answer
2018-12-21T11:07:10-0500

Qd = 200 - 2P or Pd = 100 - 0.5Q, Qs = 2P - 50 or Ps = 0.5Q - 25, a subsidy $10 per unit.

1. In equilibrium Qd = Qs, so after the imposing of a subsidy the equilibrium point is:

200 - 2P = 2(P + 10) - 50,

4P = 230,

Pe = $57.5, Qe = 200 - 2*57.5 = 85 units.

The elasticities of demand and supply in the point of equilibrium are:

Ed = -0.5*57.5/85 = -0.34,

Es = 0.5*67.5/65 = 0.34.

2. Both consumer and producer surpluses will increase.

3. Pw = 40, then consumer surplus will increase and producer surplus will decrease if the country decide to import goods, because Pw < Pe.

4. If government wants to restrict import and impose a tariff of $10, then consumer surplus will decrease, producer surplus will increase, government revenue will increase, DWL will appear.